# xFinder: Robust and Pinpoint Answer Extraction for Large Language Models

### v1

@@ 1 Introduction
Current evaluation frameworks predominantly rely on Regular Expression (RegEx) to extract answers from model outputs. For example, they may select content following phrases like "The answer is" as the key response. However, many LLMs fail to generate standardized answers [19], which complicates the extraction of key answers from their outputs. Figure 2 illustrates instances where current evaluation frameworks failed to extract responses correctly. There are two primary scenarios where regular rules fail to extract answers: first, when the model does not produce a relevant response to the evaluation question, and second, when the model’s response does not conform to standards. In the first scenario, it is unreasonable to classify a model’s failure to provide a relevant response as an error, as the evaluation aims to measure the model’s knowledge comprehension and reasoning capabilities, not merely its following ability. In the second scenario, although the LLM’s response effectively addresses the question, the response’s non-standard format impedes correct extraction using RegEx, resulting in ’false judgments’.
@@ 3 Problem Definition
    If $\boldsymbol{s}$ exists (as defined in Equation 6), $\boldsymbol{k}$ is the option within the set $\mathcal{C}$ that best corresponds to what is described by $\boldsymbol{s}$.
@@ 4 Methodology
To effectively fine-tune xFinder, a high-quality dataset is essential. To our knowledge, there is currently no dedicated dataset for Key Answer Extraction. In this section, we will detail the implementation process of xFinder, including the construction of the KAF dataset and the training of xFinder, as illustrated in Figure 3. The KAF dataset encompasses a variety of evaluation tasks, including questions, optional answer ranges, LLM responses to the questions, and the extracted key answers. This dataset is divided into three segments: a training set, a test set, and a generalization set, which are used for fine-tuning, testing, and assessing xFinder’s performance, respectively. The training set comprises 26,900 samples, the test set includes 4,961 samples, and the generalization set contains 4,482 samples. Section 4.1 details the process of LLM response collection, Section 4.2 describes our method and workflow for annotating key answers in the LLM-generated content, and Section 4.3 elaborates on our detailed approach to training xFinder.
Figure 3: Schematic of the research framework. The first three stages correspond to Sections 4.1,  4.2, and  4.3 respectively; the final stage illustrates the replacement of RegEx with xFinder in the current evaluation pipeline. The experiments in Section 5.2 demonstrate the efficacy of our approach within this pipeline. Note: The percentages 80.3%, 77.0%, and 2.2% in the center of the figure around ’unreliable evaluation’ indicate results from the Llama3-8B-Instruct model on the GSM8K benchmark using RegEx evaluation via the LM Eval Harness, OpenCompass, and UltraEval frameworks, respectively, while our method achieves a reliable result of 80.2%.
@@ 4.1 LLM Response Generation
As shown in the first part of Figure 3, we initially generate question-response pairs using different LLMs across multiple datasets. Currently, the tasks for evaluating LLM performance are quite diverse. We investigated the leading LLM benchmarks and evaluation reports to cover a wide range of popular evaluation tasks. Ultimately, we selected 18 typical evaluation task datasets, with detailed information provided in Appendix A.1. We categorized these tasks into four types, each with similar option formats, which serve as prompts to induce the LLMs to generate varied responses (as shown in the yellow boxes in Figure 2): (1) Alphabet option tasks. By referring to the design of mainstream evaluation rankings, these tasks are formatted as alphabet options, where each task aims to select the single correct option from multiple possibilities (i.e., A, B, C, D, etc.). (2) Short text option tasks. The requirement for the model to correctly map letters to option texts can affect the assessment of the model’s true capabilities [36]. Therefore, we designed tasks with key answer types as short texts in the form of alphabet options, where the goal is to select the single correct short text from multiple short texts. (3) Categorical label tasks. For traditional classification tasks (e.g., sentiment classification, topic classification, entity classification), we retained their original format, using the labels themselves as options. (4) Math option tasks. Given the output types of mathematical tasks, which could be integers, decimals, or even LaTeX code, we categorized them separately.
To further enhance the heterogeneity and richness of the dataset, we have primarily designed eight types of prompt configurations to elicit diverse responses from the LLMs. These configurations include 0-shot(-restrict444-restrict indicates that the prompt ends with the sentence ”End your final answer with ’The answer is ¡answer¿.’” to standardize the LLM’s response format.), 0-shot-cot(-restrict), 5-shot(-restrict), and 5-shot-cot(-restrict). Specific examples of these prompts are provided in Appendix E.1.
@@ 4.2 Auto Labelling and Human Recheck
The foundation of constructing the KAF dataset is the extraction of the key answer $\boldsymbol{k}$ from the LLM response $\boldsymbol{y}$, which is then used as the label. We employed distinct annotation strategies for the training, test, and generalization sets. For the training set, to improve the efficiency of the annotation process, a semi-automated procedure was adopted. This involved generating two sets of annotations from GPT-4 using varied prompts E.2. Utilizing the Self-Consistency strategy [37], items with discrepancies between the two annotation rounds, along with all mathematics questions, were consolidated for manual re-annotation using Label Studio [38]. To ensure the validity and reliability of the test and generalization sets, all labels needed to be accurate, necessitating two rounds of manual annotation for these datasets. Detailed information about the annotation methods can be found in Appendix A.3.
@@ 4.3 Training xFinder
During the training process, we primarily utilized the XTuner [39] tool developed by the InternLM team for fine-tuning. To fully exploit the instruction-following potential of xFinder, we meticulously crafted prompts as demonstrated in Appendix E.3. Additionally, to enhance the generalization capability of xFinder, we implemented data augmentation on the KAF training set, which included:
@@ 5 Experiments
We employed the QLoRA method from the XTuner framework [39] [40] for fine-tuning the foundation models, with specific tuning parameters detailed in Appendix B. To assess the efficacy of xFinder, we conducted tests using the KAF test set. We set the current mainstream frameworks—OpenCompass, LM Eval Harness, and UltraEval—with their RegEX extraction modules as baselines. The comparison metric is extraction accuracy, defined as the proportion of correctly extracted key answers relative to the dataset size. Partial results are presented in Table 1; for complete experimental results, refer to Table 6 in Appendix C.1.
@@
To further explore xFinder’s extraction capabilities on not-seen tasks, we conducted tests using the KAF generalization set. Specific results are shown in Table 2. Only two xFinder models are displayed here for brevity, but the complete experimental results can be found in Table 7 of Appendix C.1. The results indicate that xFinder maintains a high level of extraction accuracy on the generalization set. In contrast, mainstream regular expression extractors perform poorly [^5] Here, the extraction accuracy of OpenCompass, LM Eval Harness, and UltraEval has significantly decreased compared to Table 1. This is because the LLM responses in the KAF test set are generated from restricted prompts, making it easier for regular expressions to extract the key answer; however, the responses in the KAF generalization set are generated from non-restricted prompts., especially in extracting key answers for mathematical problems. Even GPT-4 struggles to extract key answers from complex and varied LLM responses accurately.
@@ 5.2 Evaluation in Real-World Scenarios
The experiments conclusively demonstrate that xFinder plays a critical role within evaluation frameworks. Next, we will use xFinder for automated evaluations in real-world evaluation tasks instead of only inspecting the extraction accuracy. To balance evaluation efficiency and quality, we calculated two metrics for all xFinder models on the KAF Generalization Set: $\Delta_{acc}$ and $\Delta_{acc}/N$ (where $N$ is the parameter size of the LLM, measured in billions). Based on these metrics, xFinder-qwen1505 (with the highest $\Delta_{acc}/N$) and xFinder-llama38it (with the highest $\Delta_{acc}$) were selected. Here, $\Delta_{acc}$ represents the difference in extraction accuracy on the KAF Generalization Set between the current xFinder and the lowest extraction accuracy among the 19 xFinder models. For specific results, refer to Table 7 in Appendix C.1.
We evaluated the performance of 10 LLMs across 14 mainstream evaluation tasks using xFinder-qwen1505 and xFinder-llama38it, along with regular extraction methods based on OpenCompass, LM Eval Harness, and UltraEval. The comprehensive scoring results are provided in Appendix C.2.
@@
Currently, the parameter settings for each evaluation framework differ significantly, including preprocessing methods for task datasets, prompt template configurations, and hyperparameter settings for the models being tested. Therefore, to compare the performance of our xFinder extraction method with RegEx methods on these frameworks, we replicated only the RegEx methods on each framework and ensured consistency in other settings. This primarily included: (1) applying uniform preprocessing to the task datasets; (2) employing a consistent zero-shot approach for all evaluation tasks, with specific prompts detailed in Appendix E.1; and (3) selecting the following LLMs for comprehensive evaluation: Phi-2, Baichuan series, Gemma series, Llama series, InternLM series, and Qwen series. Additionally, to ensure the stability of experimental results and maintain consistency in parameter settings across different LLMs, we set the temperature to 0, top_p to 0.9, and top_k to 5.
@@
Figure 4 illustrates the ranking changes of LLMs across four tasks based on evaluation results extracted from different frameworks. Additional chart results can be found in Appendix D. From the analysis, we can derive the following insights:
@@
(additional) RQ3: Is setting evaluation tasks as alphabet options reliable? Existing evaluation frameworks often aim to convert all questions into multiple-choice format with alphabet options. However, as shown in Figure 4, there are significant differences in the rankings of the CommonsenseQA task when using the alphabet option format compared to the short text format (more experimental details can be found in Appendix D.1). This suggests that alphabet options may be unreliable, and we advocate for reducing the reliance on the alphabet option format in existing evaluation frameworks.
@@ A.2 Large Language Models
We selected the models listed in Table 4 from a vast array of open-source LLMs for our experiments. Among them, the specific version of GPT-4 we used for annotating the KAF dataset is GPT4-1106-preview. The training and test sets of the KAF dataset used to generate question-answer pairs include 10 LLMs: LLaMA2-7B-Chat, Baichuan2-7B-Chat, ChatGLM3-6B, Phi-2, InternLM2-7B-Chat, Qwen1.5-4B-Chat, Qwen1.5-1.8B-Chat, Qwen1.5-0.5B-Chat, Gemma-7B-it, and Gemma-2B-it. The generalization set used to generate question-answer pairs includes 8 LLMs: LLaMA3-8B-Instruct, LLaMA2-7B-Chat, ChatGLM3-6B, Phi-2, InternLM2-1.8B-Chat, Qwen1.5-14B-Chat, Qwen1.5-MoE-A2.7B-Chat, and Gemma-2B-it.
@@
A.3 Annotation Details
@@
Appendix B Fine-Tuning Hyperparameters
@@
Appendix C Detailed Experimental Results
C.1 Extraction Accuracy Results
@@
C.2 Evaluation in Real-World Scenarios Results
C.2.1 Alphabet & Short Text Option Tasks
@@
C.2.2 Categorical Label Option Tasks

### v2

@@ 1 Introduction
Current evaluation frameworks predominantly rely on Regular Expression (RegEx) to extract answers from model outputs. For example, they may select content following phrases like "The answer is" as the key response. However, many LLMs fail to generate standardized answers [19], which complicates the extraction of key answers from their outputs. Figure 2 illustrates instances where current evaluation frameworks failed to extract responses correctly. There are two primary scenarios where regular rules fail to extract answers: first, when the model does not produce a relevant response to the evaluation question, and second, when the model’s response does not conform to standards. In the first scenario, it is unreasonable to classify a model’s failure to provide a relevant response as an error, as the evaluation aims to measure the model’s knowledge comprehension and reasoning capabilities, not merely its instruction-following ability. In the second scenario, although the LLM’s response effectively addresses the question, the response’s non-standard format impedes correct extraction using RegEx, resulting in ’false judgments’.
@@ 3 Problem Definition
    If $\boldsymbol{s}$ exists, $\boldsymbol{k}$ is the option within the set $\mathcal{C}$ that best corresponds to what is described by $\boldsymbol{s}$, as defined in Equation 6:
@@ 4 Methodology
To effectively fine-tune xFinder, a high-quality dataset is essential. To our knowledge, there is currently no dedicated dataset for Key Answer Extraction. In this section, we will detail the implementation process of xFinder, including the construction of the KAF dataset and the training of xFinder, as illustrated in Figure 3. The KAF dataset encompasses a variety of evaluation tasks, including questions, optional answer ranges, LLM responses to the questions, and the extracted key answers. This dataset is divided into three segments: a training set, a test set, and a generalization set, which are used for fine-tuning, testing, and assessing xFinder’s performance, respectively. The training set comprises 26,900 samples, the test set includes 4,961 samples, and the generalization set contains 4,482 samples. For more detailed information, please refer to the Appendix B.1. Section 4.1 details the process of LLM response collection, Section 4.2 describes our method and workflow for annotating key answers in the LLM-generated content, and Section 4.3 elaborates on our detailed approach to training xFinder.
Figure 3: Schematic of the research framework. The first three stages correspond to Sections 4.1,  4.2, and  4.3 respectively; the final stage illustrates the replacement of RegEx with xFinder in the current evaluation pipeline. The experiments in Section 5.2 demonstrate the efficacy of our approach within this pipeline. Note: The percentages 80.3%, 77.0%, and 2.2% in the center of the figure around ’unreliable evaluation’ indicate results from the Llama3-8B-Instruct on the GSM8K benchmark using RegEx evaluation via the LM Eval Harness, OpenCompass, and UltraEval frameworks, respectively, while our method achieves a reliable result of 80.2%.
@@ 4.1 LLM Response Generation
As shown in the first part of Figure 3, we initially generate question-response pairs using different LLMs across multiple datasets. Currently, the tasks for evaluating LLM performance are quite diverse. We investigated the leading LLM benchmarks and evaluation reports to cover a wide range of popular evaluation tasks. Ultimately, we selected 18 typical evaluation task datasets, with detailed information provided in Appendix A.1. We categorized these tasks into four types, each with similar option formats, which serve as prompts to induce the LLMs to generate varied responses (as shown in the left boxes in Figure 2): (1) Alphabet option tasks. By referring to the design of mainstream evaluation rankings, these tasks are formatted as alphabet options, where each task aims to select the single correct option from multiple possibilities (i.e., A, B, C, D, etc.). (2) Short text option tasks. The requirement for the model to correctly map letters to option texts can affect the assessment of the model’s true capabilities [36]. Therefore, we designed tasks with key answer types as short texts in the form of alphabet options, where the goal is to select the single correct short text from multiple short texts. (3) Categorical label tasks. For traditional classification tasks (e.g., sentiment classification, topic classification, entity classification), we retained their original format, using the labels themselves as options. (4) Math option tasks. Given the output types of mathematical tasks, which could be integers, decimals, or even LaTeX code, we categorized them separately.
To further enhance the heterogeneity and richness of the dataset, we have primarily designed eight types of prompt configurations to elicit diverse responses from the LLMs. These configurations include 0-shot(-restrict444-restrict indicates that the prompt ends with the sentence ”End your final answer with ’The answer is ¡answer¿.’” to standardize the LLM’s response format.), 0-shot-cot(-restrict), 5-shot(-restrict), and 5-shot-cot(-restrict). Specific examples of these prompts are provided in Appendix F.1.
@@ 4.2 Auto Labelling and Human Recheck
The foundation of constructing the KAF dataset is the extraction of the key answer $\boldsymbol{k}$ from the LLM response $\boldsymbol{y}$, which is then used as the label. We employed distinct annotation strategies for the training, test, and generalization sets. For the training set, to improve the efficiency of the annotation process, a semi-automated procedure was adopted. This involved generating two sets of annotations from GPT-4 using varied prompts F.2. Utilizing the Self-Consistency strategy [37], items with discrepancies between the two annotation rounds, along with all mathematics questions, were consolidated for manual re-annotation using Label Studio [38]. To ensure the validity and reliability of the test and generalization sets, all labels needed to be accurate, necessitating two rounds of manual annotation for these datasets. Detailed information about the annotation methods can be found in Appendix B.2.
@@ 4.3 Training xFinder
During the training process, we primarily utilized the XTuner [39] tool developed by the InternLM team for fine-tuning. To fully exploit the instruction-following potential of xFinder, we meticulously crafted prompts as demonstrated in Appendix F.3. Additionally, to enhance the generalization capability of xFinder, we implemented data augmentation on the KAF training set, which included:
@@ 5 Experiments
We employed the QLoRA method from the XTuner framework [39] [40] for fine-tuning the foundation models, with specific tuning parameters detailed in Appendix C. To assess the efficacy of xFinder, we conducted tests using the KAF test set. We set the current mainstream frameworks—OpenCompass, LM Eval Harness, and UltraEval—with their RegEX extraction modules as baselines. The comparison metric is extraction accuracy, defined as the proportion of correctly extracted key answers relative to the dataset size. Partial results are presented in Table 1; for complete experimental results, refer to Table 18 in Appendix D.1.
@@
To further explore xFinder’s extraction capabilities on not-seen tasks, we conducted tests using the KAF generalization set. Specific results are shown in Table 2. Only two xFinder models are displayed here for brevity, but the complete experimental results can be found in Table 19 of Appendix D.1. The results indicate that xFinder maintains a high level of extraction accuracy on the generalization set. In contrast, mainstream regular expression extractors perform poorly [^5] Here, the extraction accuracy of OpenCompass, LM Eval Harness, and UltraEval has significantly decreased compared to Table 1. This is because the LLM responses in the KAF test set are generated from restricted prompts, making it easier for regular expressions to extract the key answer; however, the responses in the KAF generalization set are generated from non-restricted prompts., especially in extracting key answers for mathematical problems. Even GPT-4 struggles to extract key answers from complex and varied LLM responses accurately.
@@ 5.2 Evaluation in Real-World Scenarios
The experiments conclusively demonstrate that xFinder plays a critical role within evaluation frameworks. Next, we will use xFinder for automated evaluations in real-world evaluation tasks instead of only inspecting the extraction accuracy. To balance evaluation efficiency and quality, we calculated two metrics for all xFinder models on the KAF Generalization Set: $\Delta_{acc}$ and $\Delta_{acc}/N$ (where $N$ is the parameter size of the LLM, measured in billions). Based on these metrics, xFinder-qwen1505 (with the highest $\Delta_{acc}/N$) and xFinder-llama38it (with the highest $\Delta_{acc}$) were selected. Here, $\Delta_{acc}$ represents the difference in extraction accuracy on the KAF Generalization Set between the current xFinder and the lowest extraction accuracy among the 19 xFinder models. For specific results, refer to Table 19 in Appendix D.1.
We evaluated the performance of 10 LLMs across 14 mainstream evaluation tasks using xFinder-qwen1505 and xFinder-llama38it, along with regular extraction methods based on OpenCompass, LM Eval Harness, and UltraEval. The comprehensive scoring results are provided in Appendix D.2.
@@
Currently, the parameter settings for each evaluation framework differ significantly, including preprocessing methods for task datasets, prompt template configurations, and hyperparameter settings for the models being tested. Therefore, to compare the performance of our xFinder extraction method with RegEx methods on these frameworks, we replicated only the RegEx methods on each framework and ensured consistency in other settings. This primarily included: (1) applying uniform preprocessing to the task datasets; (2) employing a consistent zero-shot approach for all evaluation tasks, with specific prompts detailed in Appendix F.1; and (3) selecting the following LLMs for comprehensive evaluation: Phi-2, Baichuan series, Gemma series, Llama series, InternLM series, and Qwen series. Additionally, to ensure the stability of experimental results and maintain consistency in parameter settings across different LLMs, we set the temperature to 0, top_p to 0.9, and top_k to 5.
@@
Figure 4 illustrates the ranking changes of LLMs across four tasks based on evaluation results extracted from different frameworks. Additional chart results can be found in Appendix E. From the analysis, we can derive the following insights:
@@
(additional) RQ3: Is setting evaluation tasks as alphabet options reliable? Existing evaluation frameworks often aim to convert all questions into multiple-choice format with alphabet options. However, as shown in Figure 4, there are significant differences in the rankings of the CommonsenseQA task when using the alphabet option format compared to the short text format (more experimental details can be found in Appendix E.1). This suggests that alphabet options may be unreliable, and we advocate for reducing the reliance on the alphabet option format in existing evaluation frameworks.
@@ A.2 Large Language Models
We selected the models listed in Table 4 from a vast array of open-source LLMs for our experiments.
@@
Appendix B KAF Dataset
B.1 Dataset Structure
The KAF dataset is designed to enhance the accuracy and reliability of key answer extraction in LLMs. It is divided into three segments: training set, test set, and generalization set, each serving distinct purposes in the development and evaluation of the xFinder model.
B.1.1 Training Set
The training set contains 26,900 samples. Detailed information about the models, key answer types, and prompt configurations used in the training set is presented in Tables  5,  6,  7, and  8, respectively.
LLMs Count
Baichuan2-7B-Chat 3726
Qwen1.5-4B-Chat 3051
Anonymity 2972
Qwen1.5-1.8B-Chat 2724
Gemma-7B-it 2409
InternLM2-7B-Chat 2383
Gemma-2B-it 2214
LLaMA2-7B-Chat 2102
Qwen1.5-0.5B-Chat 1927
ChatGLM3-6B 1755
PHI2 1637
Table 5: Number of samples from each LLM in the training set. Asterisk (*) denotes our manually expanded enhanced data.
Dataset Count
AgNews 1813
Amazon 2029
ARC-c (alpha) 2452
ARC-c (enh) 577
ARC-c (text) 2549
ARC-e (alpha) 2403
ARC-e (enh) 581
ARC-e (text) 2670
DBPedia 2303
GSM8K 3394
MATH 1540
MMLU (alpha) 2439
MMLU (enh) 574
MMLU (text) 1576
Table 6: Number of samples for each dataset in the training set. "enh" indicates the enhanced alphabet option questions.
Key Answer Type Count
alphabet option 9026
short text 6795
categorical label 6145
math 4934
Table 7: Number of samples for each key answer type in the training set.
Prompt Configuration Count
random-0-shot 3341
random-0-shot-cot 5201
random-0-shot-cot-restrict 2976
random-0-shot-restrict 2350
random-5-shot 3267
random-5-shot-cot 5321
random-5-shot-cot-restrict 2390
random-5-shot-restrict 2054
Table 8: Number of samples for each prompt configuration in the training set.
B.1.2 Test Set
The test set comprises 4,961 samples. Detailed information about the models, key answer types, and prompt configurations used in the test set is presented in Tables  9,  10,  11, and  12, respectively.
LLMs Count
Baichuan2-7B-Chat 786
Qwen1.5-4B-Chat 669
Qwen1.5-1.8B-Chat 545
InternLM2-7B-Chat 541
Gemma-7B-it 495
LLaMA2-7B-Chat 436
Phi-2 412
Gemma-2B-it 398
ChatGLM3-6B 344
Qwen1.5-0.5B-Chat 335
Table 9: Number of samples from each LLM in the test set.
Dataset Count
AgNews 478
Amazon 600
ARC-c (alpha) 651
ARC-c (text) 345
ARC-e (alpha) 584
ARC-e (text) 387
DBPedia 512
GSM8K 663
MMLU (alpha) 652
MMLU (text) 89
Table 10: Number of samples from each dataset in the test set.
Key Answer Type Count
alphabet option 1887
short text 821
categorical label 1590
math 663
Table 11: Number of samples for each key answer type in the test set.
Prompt Configuration Count
random-0-shot 1157
random-0-shot-cot 1248
random-5-shot 1270
random-5-shot-cot 1286
Table 12: Number of samples for each prompt configuration in the test set.
B.1.3 Generalization Set
The generalization set comprises 4,482 samples. Detailed information about the models, key answer types, and prompt configurations used in the generalization set is presented in Tables  13,  14,  15, and  16, respectively.
LLMs Count
LLaMA2-7B-Chat 562
InternLM2-1.8B-Chat 562
Gemma-2B-it 561
Qwen1.5-14B-Chat 560
Llama3-8B-Instruct 560
Qwen1.5-MoE-A2.7B-Chat 560
ChatGLM3-6B 559
Phi-2 558
Table 13: Number of samples from each LLM in the generalization set.
Dataset Count
BoolQ 320
CommonsenseQA (alpha) 320
CommonsenseQA (text) 322
hellaswag (alpha) 320
MetaMathQA 319
MultiArith 321
OpenbookQA (alpha) 320
OpenbookQA (text) 320
QNLI 320
SIQA (alpha) 320
SIQA (text) 320
Subj 320
TREC 320
WiC 320
Table 14: Number of samples from each dataset in the generalization set.
Key Answer Type Count
alphabet option 1280
short text 962
categorical label 1600
math 640
Table 15: Number of samples for each key answer type in the generalization set.
Prompt Configuration Count
random-0-shot 1122
random-0-shot-cot 1120
random-5-shot 1119
random-5-shot-cot 1121
Table 16: Number of samples for each prompt configuration in the generalization set.
B.2 Human Annotation Details
@@
Appendix C Fine-Tuning Hyperparameters
@@
Appendix D Detailed Experimental Results
D.1 Extraction Accuracy Results
@@
D.2 Evaluation in Real-World Scenarios Results
D.2.1 Alphabet & Short Text Option Tasks
@@
D.2.2 Categorical Label Option Tasks